OpenStudy (anonymous):

which of the following have property known as mass ? 1.protons,neutron,electrons 2. electrons only3. neutron & electrons 4 protons& neutrons

OpenStudy (anonymous):

Protons and neutrons have nearly the same mass while electrons are much less massive. If we assume that a neutron has a mass of 1, then the relative masses are: Neutron = 1 Proton = 0.99862349 Electron = 0.00054386734 Said another way, protons are only about 99.86% as massive as neutrons while electrons are only about 0.054% as massive as neutrons. While relative masses are nice if you want to compare protons, neutrons and electrons to one another, it doesn't tell you what the actual masses of these particles are. In kilograms, the masses are: Neutron = 1.6749286*10-27 kg Proton = 1.6726231*10-27 kg Electron = 9.1093897*10-31 kg There is another unit, called an electron volt (eV), that scientists use when talking about small things like protons, neutrons and electrons. An electron volt is actually a measurement of energy, but scientists can get away with using it to measure mass since mass and energy are related by Einstein's famous equation, E = mc2. So, in terms of MeV (Megaelectron volts, 1 MeV = 1,000,000 eV), the masses are: Neutron = 939.56563 MeV Proton = 938.27231 MeV Electron = 0.51099906 MeV Author: Steve Gagnon, Science Education Specialist (Other answers by Steve Gagnon) Related Pages: If different types of quarks have different masses, then why are protons and neutrons said to have the same mass?

OpenStudy (anonymous):

so ANSWER IS 1

OpenStudy (anonymous):

@sraone88 thanks so much!